There are three 20.0 ohm resistors in series across a 120.0-V generator. What is the equivalent resistance of the circuit?
What is the current in the circuit?
What is the total voltage drop across the circuit?
1
Expert's answer
2018-04-29T14:02:08-0400
Equivalent resistance of the circuit:
R_e=R_1+R_2+R_3=20+20+20=60 Ohm
The current in the circuit:
I=U/R_e =120/60=2 A
Total voltage drop across the circuit:
U_1=I·R_1=2·20=40 V
U_2=I·R_2=2·20=40 V
U_3=I·R_3=2·20=40 V
U_∑=U_1+U_2+U_3=40+40+40=120 V